OpenStudy (anonymous):

metals will be given 1 question Frank has a circular garden. The area of the garden is 100 ft2. What is the approximate distance from the edge of Frank’s garden to the center of the garden? (A = πr2) 1.6ft 2.8ft 3.10ft 4. 12ft

OpenStudy (anonymous):

solve \[100=\pi r^2\] for \(r\)

OpenStudy (anonymous):

divide by \(\pi\) then take the square root use a calculator

OpenStudy (anonymous):

about \(5.64\) so pick "6"

OpenStudy (anonymous):

ha ha metals!!!!!!! anyways no probs.... r = √(100/π) = 5.64 rounding off makes it 6 ft

OpenStudy (anonymous):

omggggg thank you so much i went on wolffram but id idnt know how to use it but thanks @satellite73

OpenStudy (anonymous):

thanks too @ryanvarghese12790

OpenStudy (anonymous):

yw